# Efficient Bayesian Additive Regression Models For Microbiome Studies

### v1

@@ Abstract
Statistical analysis of microbiome data is challenging. Bayesian multinomial logistic-normal (MLN) models have gained popularity due to their ability to account for the count compositional nature of these data. However, these models are often computationally intractable to infer. Recently, we developed a computationally efficient and accurate approach to inferring MLN models with a Marginally Latent Matrix-T Process (MLTP) form: MLN-MLTPs. Our approach is based on a novel sampler with a marginal Laplace approximation – called the Collapse-Uncollapse (CU) sampler. However, existing work with MLTPs has been limited to linear models or models of a single non-linear process. Moreover, existing methods lack an efficient means of estimating model hyperparameters. This article addresses both deficiencies. We introduce a new class of MLN Additive Gaussian Process models (MultiAddGPs) for deconvolution of overlapping linear and non-linear processes. We show that MultiAddGPs are examples of MLN-MLTPs and derive an efficient CU sampler for this model class. Moreover, we derive efficient Maximum Marginal Likelihood estimation for hyperparameters in MLTP models by taking advantage of Laplace approximation in the CU sampler. We demonstrate our approach using simulated and real data studies. Our models produce novel biological insights from a previously published artificial gut study.
Additive Regression,
Multinomial Logistic-Normal Models,
Bayesian Inference,
Microbiome Data,
keywords:
[…]
, and
@@ 1 Introduction
Dysregulation of human-, animal-, and even plant-associated microbial communities (microbiota) are known to cause disease (Ballen et al., 2016; Frati et al., 2018; Holleran et al., 2018; Sharon et al., 2019; Gao et al., 2021). In humans, alterations of microbiota play a causal role in obesity (Tilg et al., 2011; Ley, 2010), inflammatory bowel disease (Honda and Littman, 2012; Glassner, Abraham and Quigley, 2020; Kostic, Xavier and Gevers, 2014), and even cancer (Schwabe and Jobin, 2013; Helmink et al., 2019). As a result, many researchers study how dietary, host physiologic, and environmental factors influence the relative abundance of different bacterial taxa in microbiota. These factors can have linear or non-linear effects on community structure (Cheng et al., 2019; Sankaran and Jeganathan, 2024; Schwager et al., 2017). Even within closed in vitro systems, microbiota can demonstrate non-linear temporal variation (Silverman et al., 2018). Overall, flexible statistical methods are needed to disentangle linear and non-linear effects on microbiota.
@@
Beyond the biological complexity of microbiota, limitations of the measurement process further complicate analyses. These data are typically represented as a $D\times N$ count table $Y$ with elements $Y_{dn}$ denoting the number of DNA molecules from taxon $d$ observed (sequenced) in sample $n$. Critically, the counts $Y_{dn}$ do not represent the true abundance of taxon $d$ in the biological system (e.g., colon) from which the sample $n$ was obtained. Instead, the counts are the results of a random sample of that pool of microbes. Due to limitations of the measurement process, the size of that sample (the sequencing depth; $\sum_{d=1}^{D}Y_{dn}$) is typically arbitrary and unrelated to the total microbial load in the system (Vandeputte et al., 2017). As a result, many authors call these data compositional, reflecting the idea that the data only provide information about the relative abundances of the different taxa within each community (Gloor et al., 2017; McGregor et al., 2023; Mao and Ma, 2022). However unlike traditional compositional data (e.g., continuous simplex valued data), these data are zero-laden counts. The presence of counting noise and zeros limits standard approaches to compositional data analysis, which typically involve analyses of log-ratio transformed data (Kuczynski et al., 2012; Silverman et al., 2018, 2017, 2020; Kaul et al., 2017; Li, 2015; Gloor et al., 2016).
Many researchers have turned to Bayesian Multinomial Logistic Normal (MLN) models to address the measurement process’s challenges (Silverman et al., 2021; Äijö, Müller and Bonneau, 2018; Grantham et al., 2020; Silverman et al., 2018, 2022). The multinomial is used to model uncertainty due to random counting, while the logistic normal captures the extra-multinomial variability typically seen in these data (Silverman et al., 2022). More conceptually, Bayesian MLN models allow researchers to model community composition as a latent simplicial vector that is informed by the observed count data. Moreover, unlike the more well-known Dirichlet distribution, the logistic-normal has a rich covariance structure which allows modeling both positive and negative covariation between taxa (Aitchison and Shen, 1980; Silverman et al., 2018). The logistic-normal is also self-conjugate (as it is multivariate normal under a suitable log-ratio transformation), allowing for a wide variety of models to be built in the latent simplex space. However, the multinomial and the logistic-normal are not conjugate, making inference of these models computationally challenging or even intractable.
Recent advances have made Bayesian MLN models practical for microbiota analyses. Early Bayesian MLN models were inferred via Metropolis-within-Gibbs and only scaled to a handful of dimensions (Cargnoni, Müller and West, 1997). Pólya-Gamma data augmentation has been used successfully in some instances (Glynn et al., 2019); however, it is still too computationally intensive for widespread use due to the inability to perform blocked Gibbs updates of Pólya-Gamma random variables while maintaining the logistic-normal form (Linderman, Johnson and Adams, 2015). Subsequently, researchers successfully used Hamiltonian Monte Carlo (Äijö, Müller and Bonneau, 2018; Grantham et al., 2020; Silverman et al., 2018). Yet scalability still limits those methods. For example, the sampler used in Silverman et al. (2018) took more than four hours to analyze a dataset of approximately one thousand samples, yet only ten taxa. More recently, we proved that a wide variety of Bayesian MLN models, including generalized linear models and generalized Gaussian process regression models, share a common marginal form called a Latent Matrix-T Process (LTP) (Silverman et al., 2022). We showed that a Laplace approximation to this marginal form was extremely accurate, leading to an efficient and accurate approximate inference procedure called the Collapse-Uncollapse (CU) sampler. Our result demonstrated that this approach is often 4-5 orders of magnitude faster than HMC-based methods with minimal error in posterior calculations (Silverman et al., 2022).
Despite these advances, there remains a dearth of tools for disentangling the effects of multiple measured factors on microbiota. Recently, Cheng et al. (2019) proposed an additive Gaussian process framework to address this need. Yet their approach assumed the data was transformed Gaussian, ignoring count compositional nature of these data. Moreover, our prior work with MLTPs was limited to factors that have a linear effect on microbial composition (generalized linear models) or a single factor that had a nonlinear effect (generalized Gaussian process regression models). Finally, our prior work did not address the problem of hyperparameter selection, which is often critical in both linear and nonlinear modeling (e.g., selection of kernel parameters).
This article addresses the limitations of prior methods and develops a flexible, and computationally efficient approach to disentangling both linear and nonlinear effects on microbiota. As in Cheng et al. (2019), our approach is based on a class of additive Gaussian process regression models. Unlike Cheng et al. (2019), we do not assume that the data is transformed Gaussian and instead prove that Bayesian MLN Additive Gaussian Process Models are also MLTPs. Using those results, we extend the CU sampler to this class of models. Moreover, we extend the results of Silverman et al. (2022) by developing an efficient maximum marginal likelihood estimator for model hyperparameters. Through simulation studies, we validate our methods; we find posterior intervals from our models often cover the truth ten to a hundred times more often than those of transformed Gaussian models. Finally, we apply our models to a previously published longitudinal in vitro gut microbiome study which studied the effects of starvation on microbial communities (Silverman et al., 2018). Our models are able to provide the first estimates of the impact of starvation over time within that study.
We organize this article as follows. Section 2 provides a review of the Marginally LTP class of Bayesian MLN models, as well as a review of the Collapse-Uncollapse sampler with marginal Laplace approximation. Section 3 presents a multinomial logistic-normal generalized additive Gaussian process regression (MultiAddGP) model and proves that this model is part of the Marginally LTP class. We also derive a computationally efficient approach to hyperparameter estimation in MLTP models via maximum marginal likelihood. Sections 4 and 5 demonstrate our approach through application to both simulated and real microbiome data. Finally, we conclude with a discussion in Section 6.
@@
2 Review of Marginally Latent Matrix-T Processes and the CU Sampler
We describe the class of Marginally Latent Matrix-T Processes (MLTPs) by sequentially generalizing from Matrix-T Processes to Latent Matrix-T Processes (LTPs) and finally MLTPs. We then describe the subset of Bayesian Multinomial Logistic-Normal MLTPs (MLN-MLTPs) before reviewing the inference of this class of models.
2.1 Defining Marginally Latent Matrix-T Processes (MLTPs)
@@
Just as Gaussian processes can be defined based on the marginal properties of the multivariate normal, matrix normal processes and Matrix-T processes can be defined by the marginal properties of the matrix normal and matrix-T distributions (Silverman et al., 2022). Matrix-T processes generalize Student-T processes and Gaussian processes (Silverman et al., 2022).
@@
A stochastic process $\mathbf{Y}\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A})$ defined on the set $\mathcal{W}=\mathcal{W}^{(1)}\times\mathcal{W}^{(2)}$ is a matrix-T process if $\mathbf{Y}$ evaluated on any two finite subsets $\mathcal{X}^{(1)}\subset\mathcal{W}^{(1)}$ and $\mathcal{X}^{(2)}\subset\mathcal{W}^{(2)}$ is a random matrix $\mathbf{Y}$ of dimension $|\mathcal{X}^{(1)}|\times|\mathcal{X}^{(2)}|$ that follows a matrix-T distribution: $\mathbf{Y}\sim T(\nu,\mathbf{M,V,A})$. $\nu$ is a scalar value strictly greater than zero. Let $x_{i}^{(1)},x_{j}^{(1)}\in\mathcal{X}^{(1)}$ and $x_{i}^{(2)},x_{j}^{(2)}\in\mathcal{X}^{(2)}$. $\mathbf{M}_{ij}=\mathbf{M}(x_{i}^{(1)},x_{j}^{(2)})$ is the matrix function representing the mean, and $\mathbf{V}_{ij}=\mathbf{V}(x_{i}^{(1)},x_{j}^{(1)})$ and $\mathbf{A}_{ij}=\mathbf{A}(x_{i}^{(2)},x_{j}^{(2)})$ are kernel functions.
Latent Matrix-T Processes (LTP) generalize Matrix-T processes. $\mathbf{Y}$ is said to be an LTP if
$\mathbf{Y}$ $\sim g(\mathbf{\Pi},\lambda)$

### v2

# Scalable Bayesian Semiparametric Additive Regression Models For Microbiome Studies

@@ Abstract
Statistical analysis of microbiome data is challenging. Bayesian multinomial logistic-normal (MLN) models have gained popularity due to their ability to account for the count compositional nature of these data, but existing approaches are either computationally intractable or restricted to purely parametric or non-parametric methods, which limit their flexibility and scalability. In this work, we introduce MultiAddGPs, a novel semi-parametric framework that integrates additive Gaussian Process (GP) regression within a Bayesian MLN model to disentangle linear and non-linear covariate effects, including non-stationary dynamics. Our approach builds on the computationally efficient Collapse-Uncollapse (CU) sampler and additive GP regression, introducing a novel back-sampling algorithm and marginal likelihood approximation for efficient inference and hyperparameter estimation. Our models are over 240,000 times faster than alternatives while simultaneously producing more accurate posterior estimates. Additionally, we incorporate non-stationary kernel functions designed to model treatment interventions and disease effects. We demonstrate our approach using simulated and real data studies and produce novel biological insights from a previously published human gut microbiome study. Our methods are publicly available as part of the fido software package on CRAN  [^1] A preliminary version of this work was presented at the NeurIPS 2024 Workshop on “Bayesian Decision-making and Uncertainty,” but was not published in the workshop proceedings..
Keywords: Additive Regression, Multinomial Logistic-Normal Models, Bayesian Inference, Microbiome Data, Gaussian Process
@@ 1 Introduction
Dysregulation of human-, animal-, and even plant-associated microbial communities (microbiota) are known to cause disease (Ballen et al. 2016, Frati et al. 2018, Holleran et al. 2018, Sharon et al. 2019, Gao et al. 2021). In humans, alterations of microbiota play a causal role in obesity (Tilg et al. 2011, Ley 2010), inflammatory bowel disease (Honda & Littman 2012, Glassner et al. 2020, Kostic et al. 2014), and even cancer (Schwabe & Jobin 2013, Helmink et al. 2019). As a result, many researchers study how dietary, host physiologic, and environmental factors influence the relative abundance of different bacterial taxa in microbiota. These factors can have linear or non-linear effects on community structure (Cheng et al. 2019, Sankaran & Jeganathan 2024, Schwager et al. 2017). Even within closed in vitro systems, microbiota can demonstrate non-linear temporal variation (Silverman et al. 2018). While additive regression models have been proposed to model these complexities, existing implementations are too computationally intensive for widespread use and ignore the count compositional nature of these data (Timonen et al. 2021, Cheng et al. 2019). Overall, interpretable and flexible statistical methods are needed to disentangle and infer linear and non-linear effects on microbiota.
@@
Limitations of the sequence counting measurement process complicate the analysis of microbiome data. These data are typically represented as a $D\times N$ count table $Y$ with elements $Y_{dn}$ denoting the number of DNA molecules from taxon $d$ observed (sequenced) in sample $n$. Critically, the counts $Y_{dn}$ do not represent the true abundance of taxon $d$ in the biological system (e.g., colon) from which the sample $n$ was obtained. Instead, the counts are the results of a random sample of that pool of microbes. Due to limitations of the measurement process, the size of that sample (the sequencing depth; $\sum_{d=1}^{D}Y_{dn}$) is typically arbitrary and unrelated to the total microbial load in the system (Vandeputte et al. 2017). As a result, many authors call these data compositional, reflecting the idea that the data only provide information about the relative abundances of the different taxa within each community (Gloor et al. 2017, McGregor et al. 2023, Mao & Ma 2022). However, unlike traditional compositional data (e.g., continuous simplex valued data), these data are zero-laden counts. The presence of counting noise and zeros limits standard approaches to compositional data analysis, which typically involve analyses of log-ratio transformed data (Kuczynski et al. 2012, Silverman et al. 2018, 2017, 2020, Kaul et al. 2017, Li 2015, Gloor et al. 2016).
A growing number of researchers are adopting Bayesian Multinomial Logistic Normal (MLN) models to tackle the challenges associated with the measurement process (Silverman et al. 2021, Äijö et al. 2018, Grantham et al. 2020, Silverman et al. 2018, 2022, Saxena et al. 2024). The multinomial component addresses uncertainty stemming from random counts, while the logistic normal component accounts for the extra-multinomial variability often observed in these datasets (Silverman et al. 2022). Furthermore, the multinomial framework provides a more intuitive representation of the generating process behind the numerous zero counts than traditional univariate zero-inflated models (Silverman et al. 2020). Conceptually, Bayesian MLN models allow researchers to represent community composition as a latent simplicial vector informed by the observed count data. In contrast to the more commonly used Dirichlet distribution, the logistic-normal distribution offers a rich covariance structure that facilitates the modeling of both positive and negative covariation among taxa (Aitchison & Shen 1980, Silverman et al. 2018). The logistic-normal distribution is self-conjugate (as it forms a multivariate normal distribution under a suitable log-ratio transformation), permitting the construction of a diverse range of models within the latent simplex space. Although Multinomial-logistic normals have historically posed computational challenges that hinder widespread application, recent innovations, such as the Collapse-Uncollapse (CU) Sampler (Silverman et al. 2022), have addressed this limitation. Nevertheless, there is still a need for more flexible approaches that can effectively disentangle linear and non-linear effects as existing methods do not allow for additive linear and non-linear modeling (Silverman et al. 2022).
In this study, we introduce MultiAddGPs, a versatile and interpretable semi-parametric framework aimed at addressing key challenges in microbiome data analysis. MultiAddGPs are scalable Bayesian MLN models that facilitate additive combinations of both linear and non-linear Gaussian process components. Our approach demonstrates exceptional computational efficiency and scalability, outperforming existing alternatives by a factor of over 400 while yielding more accurate posterior estimates. Our inference algorithms leverage the CU Sampler from Silverman et al. (2022), while simultaneously enhancing those methods. We introduce a novel back-sampling algorithm for efficient posterior simulation from multivariate additive Gaussian process models as well as non-stationary kernels for estimation of non-linear intervention effects on complex microbial communities.
Moreover, we introduce an accurate and efficient method for marginal likelihood estimation that leverages the Laplace approximation within the CU sampler. Our marginal likelihood estimation procedure is critical as it allow for efficient hyperparameter estimation for MultiAddGPs as well as the broader class of models that can be sampled by the CU sampler (Silverman et al. 2022). We validate our methods through simulation and real data studies. We demonstrate that MultiAddGPs can directly estimate complex, time-lagged effects of interventions on human microbiota. In an application involving an ex vivo artificial human gut study, our findings offer the first evidence that microbiota communities display oscillatory-like behavior in response to short-term dietary interventions.
MultiAddGPs MLTPs Lgpr LongGP
Model Linear and Nonlinear Linear or Nonlinear Nonlinear Nonlinear
Flexible Kernels
Scalable
Hyperparameter Estimation
Count Compositional
Table 1: A comparison of MultiAddGPs and existing alternatives. MultiAddGPs (this article) facilitate additive modeling of both linear and non-linear processes. In contrast, MLTPs (Silverman et al. 2022) consist of either a single linear or a single non-linear component. The Lgpr (Timonen et al. 2021) and LongGP (Cheng et al. 2019) frameworks permit only additive non-linear components. Both MultiAddGPs and MLTPs are implemented in the fido software package available on CRAN, offering both pre-defined and user-defined kernels. Conversely, the Lgpr and LongGP packages restrict usage to pre-defined kernels. MultiAddGPs and MLTPs are computationally efficient due to the CU Sampler, while LongGP and Lgpr are computationally intensive and do not scale effectively to the datasets analyzed in this article. Notably, the MLTPs proposed by Silverman et al. (2022) lack a mechanism for estimating model hyperparameters, a feature present in the other three methods. Furthermore, MultiAddGPs and MLTPs address the count compositional nature of the data using Bayesian MLN models, whereas Lgpr is limited to univariate likelihoods, and LongGP assumes the data follows a Gaussian distribution.
@@
2 Methods
2.1 Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGPs)
Let $\mathbf{Y}_{\cdot n}$ denote a $D$-vector of observed data, $\mathbf{X}_{\cdot n}$ a $Q_{0}$-vector of covariates to model linearly, and each $\mathbf{Z}^{(k\in\{1,\dots,K\})}_{\cdot n}$ a $Q_{k}$-vector of covariates to be modeled with distinct non-linear functions. The MultiAddGPs model have the following form:
$\boldsymbol{\Pi}_{\cdot n}$ $=\phi^{-1}(\mathbf{H}_{\cdot n})$ (2)
$\mathbf{H}_{\cdot n}$ $\sim\mathcal{N}(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ (3)
with priors $\mathbf{B}\sim\mathcal{N}(\boldsymbol{\Theta}^{(0)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(0)})$, $\mathbf{f}^{(k)}\sim\mathcal{GP}(\boldsymbol{\Theta}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(k)})$, and $\boldsymbol{\Sigma}\sim\mathcal{IW}(\boldsymbol{\Xi},\zeta)$. Here, the likelihood is $p(\mathbf{Y}_{\cdot 1},\dots,\mathbf{Y}_{\cdot N})\sim\prod_{n=1}^{N}\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$, $\boldsymbol{\Sigma}$ is a $(D-1)\times(D-1)$ covariance matrix, and the transform $\phi$ is an invertible log-ratio transform from the $D$-dimensional simplex to $D-1$ dimentional real-space: $\mathbf{H}_{\cdot n}=\phi(\boldsymbol{\Pi}_{\cdot n}\in\mathbb{S}^{D})\in\mathbb{R}^{D-1}$. For computational efficiency, we used the following Additive Log-Ratio (ALR) transform, which takes the $D$-th taxa as a reference:
$\mathbf{H}_{.n}=\phi(\boldsymbol{\Pi}_{.n})=\left\{\log\left(\frac{\pi_{1n}}{\pi_{Dn}}\right),\ldots,\log\left(\frac{\pi_{(D-1)n}}{\pi_{Dn}}\right)\right\}^{T}.$
There is no loss of generality as posterior samples taken with respect to the $\text{ALR}_{D}$ coordinate system can be transformed into any other log-ratio coordinate system (Pawlowsky-Glahn et al. 2015, Appendix A.3). For context, the $\text{ALR}_{D}$ transform is the inverse of the identified softmax transform.
For the matrix-normal prior on the linear term, $\boldsymbol{\Theta}^{(0)}$ is the mean matrix and $\boldsymbol{\Gamma}^{(0)}$ is a $Q_{0}\times Q_{0}$ covariance matrix representing covariance in the parameters of the $Q_{0}$ covariates. The terms $\boldsymbol{\Theta}^{(k)}$ and $\boldsymbol{\Gamma}^{(k)}$ in the $K$ matrix-normal process priors echo their linear counterparts but are functions (e.g., mean and kernel functions) rather than fixed dimensional matrices. Additionally, we permit hyperparameters $\Omega$ in any part of the model, including but not limited to hyperparameters in the kernel functions (e.g., bandwidth parameters) or hyperparameters in the mean functions.
As we will show through simulated and real data analyses in Sections 3 and 4, this is a very flexible class of models which can be used in a wide range of additive linear and non-linear modeling tasks.
2.2 Posterior Estimation in MultiAddGPs
We develop scalable inference for MultiAddGPs using a two-step Collapse-Uncollapse (CU) sampler (Silverman et al. 2022). Any multi-parameter Bayesian model $p(\mathbf{H},\mathbf{\Phi}\mid\mathbf{Y})$ can be sampled in two-step: (1) obtain a sample from the marginal $p(\mathbf{H}\mid\mathbf{Y})$ (the collapse step); (2) conditioned on that sample, obtain a sample from the corresponding conditional $p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})$ (the uncollapse step). Yet this approach is often infeasible as these posterior marginals and conditionals are rarely available for non-conjugate models. Full details on the CU sampler can be found in Supplementary Section 1. Silverman et al. (2022) showed that there was a large class of semi-conjugate models called Marginally Latent matrix-T Processes (MLTPs) where, the posterior conditional $p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})$ had a closed form and the posterior marginal $p(\mathbf{H}\mid\mathbf{Y})$ was the posterior of a Latent matrix-T Process (LTP) which could be accurately and efficiently sampled through a Laplace approximation. Below we show that MultiAddGPs are MLTPs and derive the parameters of their marginal LTP form. These results, are used to form the Laplace approximation to the Collapse step for MultiAddGPs. We then derive a novel Back-sampling (BS) for multivariate additive Gaussian process regression models. The BS algorithm provides a closed-form solution for the Uncollapse step for MultiAddGPs. Finally, we develop efficient hyperparameter inference for MLTPs (including MultiAddGPs) by maximum marginal likelihood estimation and discuss the identification of posterior samples from MultiAddGPs.
Algorithm 1 The Collapse-Uncollapse (CU) Sampler for MultiAddGPs Models
Input: Data $\{\mathbf{Y},\mathbf{X},\mathbf{Z}^{(1)},\dots,\mathbf{Z}^{(K)}\}$ [^2] We omit conditioning on covariates $\mathbf{X}$ and $\mathbf{Z}$ in the notation for brevity., and prior parameters $\Delta=\{\mathbf{\Theta}^{(0)},\dots,\mathbf{\Theta}^{(K)},\mathbf{\Gamma}^{(0)},\dots,\mathbf{\Gamma}^{(K)},\boldsymbol{\Xi},\zeta\}$.
Output: $S$ samples of $\{\mathbf{H},\boldsymbol{\Sigma},\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\}$.
1:for each sample $s\in\{1,\dots,S\}$ do
2:     Sample $\mathbf{H}^{[s]}\sim p(\mathbf{H}|\mathbf{Y},\Delta)$ where $p(\mathbf{H},\mathbf{Y},\Delta)$ is an LTP; $\triangleright$ Collapsed Step
3:     Sample $\{\boldsymbol{\Sigma}^{[s]},\mathbf{F}^{[s]}\}\sim p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H}^{[s]},\Delta)$; $\triangleright$ Uncollapsed Step (1)
4:     Sample $\mathbf{B}^{[s]},\mathbf{f}^{(1)[s]},\dots,\mathbf{f}^{(K)[s]}=\text{Back-Sampling}(\mathbf{Y},\mathbf{F}^{[s]},\boldsymbol{\Sigma}^{[s]},\Delta)$; $\triangleright$ Uncollapsed Step (2)
5:end for
2.2.1 MultiAddGPs as MLTPs and the Collapse Step
@@
Just as Gaussian processes can be defined based on the marginal properties of the multivariate normal, matrix normal processes and matrix-t processes can be defined by the marginal properties of the matrix normal and matrix-t distributions (Silverman et al. 2022). Matrix-t processes generalize Student-t processes and Gaussian processes (Silverman et al. 2022).
@@
A stochastic process $\mathbf{Y}\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A})$ defined on the set $\mathcal{W}=\mathcal{W}^{(1)}\times\mathcal{W}^{(2)}$ is a matrix-T process if $\mathbf{Y}$ evaluated on any two finite subsets $\mathcal{X}^{(1)}\subset\mathcal{W}^{(1)}$ and $\mathcal{X}^{(2)}\subset\mathcal{W}^{(2)}$ is a random matrix of dimension $|\mathcal{X}^{(1)}|\times|\mathcal{X}^{(2)}|$ that follows a matrix-T distribution: $\mathbf{Y}\sim T(\nu,\mathbf{M,V,A})$. $\nu$ is a scalar value strictly greater than zero. Let $x_{i}^{(1)},x_{j}^{(1)}\in\mathcal{X}^{(1)}$ and $x_{i}^{(2)},x_{j}^{(2)}\in\mathcal{X}^{(2)}$. $\mathbf{M}_{ij}=\mathbf{M}(x_{i}^{(1)},x_{j}^{(2)})$ is the matrix function representing the mean, and $\mathbf{V}_{ij}=\mathbf{V}(x_{i}^{(1)},x_{j}^{(1)})$ and $\mathbf{A}_{ij}=\mathbf{A}(x_{i}^{(2)},x_{j}^{(2)})$ are kernel functions.
Latent Matrix-T Processes (LTP) generalize Matrix-T processes.
Definition 2.2 (Latent Matrix-T Processes).
$\mathbf{Y}$ is said to be an LTP if
$\mathbf{Y}$ $\sim g(\mathbf{\Pi})$